# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v4

@@ 4.2 Dynamic Single-vision Setup
Sensor Design. The second approach offers a solution for iMarker detection using a single camera equipped with a switchable (dynamic) circular polarizer alternating between left- and right-handed polarization (Fig. 5-a), which will be described in detail in a forthcoming paper. In brief, it comprises a fixed circular polarizer paired with a nematic liquid crystal film that functions as a $\lambda/2$-plate in its relaxed (OFF) state and loses its optical functionality when an electric field is applied (ON). A $\lambda/2$-plate converts one circular polarization to its opposite; therefore, by switching the liquid crystal film between its ON and OFF states (controlled by a function generator synchronized with the camera’s frame rate), the camera sequentially captures frames with right- and left-handed polarization, separated in time. In an ideal setup, camera frame $f_{t}$ captures the scene with a particular polarization (left- or right-handed), and frame $f_{t+1}$ captures it with the opposite polarization. Thus, the function generator should send signal $\mathbf{S}_{g}$ to trigger the liquid crystal film at time $t$ based on Equation 1:
@@
The liquid crystal cell’s OFF-to-ON switching time is negligible and can be ignored; however, the ON-to-OFF relaxation is slower. Thus, the intermediate frames with undefined polarization for high frame rates may need to be discarded for the iMarker detection procedure. Once optimized for the liquid crystal cell, it ensures that CSR-coated regions in the iMarkers are blocked in one processed frame and visible in the next.
Detection Algorithm. Inspired by the “spatial” frame processing in the dual-vision setup, this approach employs a frame-level algorithm that uses “temporal” subtraction to detect the iMarker patterns. Assuming the liquid crystal film operates sufficiently fast relative to the video frame rate to produce sequential frames with alternating polarization states (as illustrated in Equation 1), the consecutive (temporal) frame subtraction algorithm to detect iMarkers using a switchable polarizer is shown in Algorithm 2. In this process, each frame $f_{t-1}^{\text{~off}}$ and its subsequent frame $f_{t}^{\text{~on}}$, captured with alternating polarization states, are subtracted to produce a grayscale image $f_{p}$, which can then be enhanced through post-processing to facilitate iMarker detection. Additionally, the current frame is stored for subtraction from its succeeding frame in the next iteration. As with the previous approaches, the final stage involves detecting fiducial markers $\mathbf{M}$ within the processed frame $f_{p}$.
@@ 4.3 Static Single-vision Setup
Sensor Design. Another configuration offers a solution for iMarker detection using a single camera equipped with a static circular polarizer attached to its lens (Fig. 5-b). The polarizer is selected to block the circular polarization reflected by the CSRs forming the iMarker pattern. For instance, a left-handed circular polarizer will block iMarker regions with CSRs reflecting right-handed polarization, and vice versa. Thus, a visible-range iMarker with CSRs reflecting right-hand polarized light on a same-color background (e.g., green, blue, etc.), indistinguishable to the human eye or standard cameras, becomes detectable as the sensor’s left-handed polarizer blocks the CSR regions, revealing the camouflaged iMarker patterns. This design is well-suited for detecting visible-range iMarkers camouflaged within their environment, while it can also be used for IR/UV-range iMarkers using proper cameras.
Detection Algorithm. Unlike the dual-vision solution, the algorithm for the static single-vision setup does not eliminate the background. However, it simplifies constructing and integrating an iMarker detection sensor by shifting more computation to the detector methodology, achieving a practical balance between complexity and performance. For the visible-range camouflaged iMarkers, the procedure is shown in Algorithm 3. Accordingly, detecting iMarkers $\mathbf{M}$ in a frame $f\in\mathbf{F}$ involves converting $f$ from RGB to HSV color space to isolate the color range corresponding to the marker. The resulting mask frame $f_{p}$ highlights the target color range $c_{r}$ as black pixels, with all other areas appearing as white. With the assistance of the polarizer, the mask frame $f_{p}$ highlights the blocked CSR regions of an iMarker with color $c$ (e.g., green) positioned against a background of the same color. Subsequent stages of the algorithm apply post-processing to eliminate noise, followed by detecting marker patterns $\mathbf{M}$ within $f_{p}$.
Detecting UV/IR-range iMarkers follows the same stages as the visible-range detection, provided that a UV/IR-range camera is used. However, color masking is unnecessary since the input is grayscale, and range thresholding is applied to identify the intensity variations in CSR regions, highlighting the iMarker patterns. Thus, Algorithm 4 can be employed for UV/IR-range iMarker detection scenarios.
@@ 5 Benchmarking and Evaluation
This section explores the potential and applicability of iMarkers from various aspects of real-world scenarios.
@@ 5.1 Operational Implementation
Considering the iMarker detection strategies outlined in Section 4, real-world sensor employment and algorithm implementation are the focus of this subsection to form the basis for subsequent evaluations.
@@
Sensor. Fig. 6 depicts the different sensor configurations developed for iMarker detection. The dual-vision setup shown in Fig. 6-a comprises two ELP 8MP HD camera modules with $75^{\circ}$ FoV facing a Thorlabs polarizing cube beamsplitter with a 50:50 split ratio, all mounted in a compact cage assembly. The other dual-vision setup in Fig. 6-b contains two synchronized U3-3271LE iDS cameras aligned with a Thorlabs polarizing plate beamsplitter housed within a custom-designed 3D-printed cage. The static single-vision setup shown in Fig. 6-c is an Intel® RealSense™ D435 camera equipped with a left-handed circular polarizer affixed to its RGB lens, effectively blocking the right-handed CSR-coated areas of camouflaged iMarkers. Finally, the dynamic single-vision sensor setup in Fig. 6-d incorporates a switchable polarizer controlled by a function generator mounted on an iPhone 13 case and aligned with its camera. While this prototype relies on manual polarizer switching, we are investigating a real-time version where the detection software dynamically synchronizes the camera and function generator to adjust polarization (left- or right-handed).
@@ 5.2 Qualitative Evaluation
Given the complexity of measuring aesthetic perception, this section aims to visually assess iMarkers and showcase their detection performance. To achieve this, various iMarkers were strategically placed in rigid real-world locations and captured using different hardware sensors, highlighting their seamless integration into diverse settings. It should be noted that while aesthetic perception is inherently subjective, the visual comparisons provided in Fig. 1 and Fig. 3 offer practical, real-world evidence supporting the unobtrusiveness of iMarkers in diverse environments.
@@
Fig. 8 depicts several qualitative results, showing the versatility of iMarkers across different spectral ranges and detection setups. Accordingly, Fig. 8(a) presents a visible-range green iMarker discreetly embedded on a similar color background, demonstrating its visual camouflage and unobtrusiveness. In this scenario, the static single-vision setup was employed to capture the scene, where the setup’s polarizer effectively blocks the CSR regions forming the inner iMarker pattern, enhancing the contrast and enabling enhanced separation for detection. The detection and pose estimation outcome was obtained using Algorithm 3. Fig. 8(b) and Fig. 8(c) depict the output of the dynamic single-vision setups applied to the same visible-range iMarker in indoor and outdoor conditions. Detection is achieved through temporal subtraction between frames $f_{t-1}^{\text{~off}}$ and $f_{t}^{\text{~on}}$ following the procedure described in Algorithm 2. In Fig. 8(d), the output of the dual-vision setup targeting a visible-range iMarker placed on a patterned surface under low-light conditions is shown. In this configuration, both cameras simultaneously capture the same scene, with the left camera acquiring frame $f_{l}$ through a left-handed polarizer and the right camera capturing frame $f_{r}$ through a right-handed polarizer. Marker detection was performed using Algorithm 1 , which applies spatial subtraction between the polarized image pairs. Finally, the UV-range iMarker placed on a transparent surface depicts exceptional invisibility to the naked eye, as illustrated in Fig. 8(e). To reveal the inner patterns of the iMarker, a near-UV illumination was applied, and the same scene was captured using the UV-range camera, as shown in the middle sub-figure. The methodology introduced in Algorithm 4 facilitates the marker detection process by applying a tailored thresholding strategy to the UV image, enabling reliable iMarker’s invisible patterns extraction.
@@ 5.3 Detection Range and Pose Accuracy Evaluation
To evaluate the effectiveness of iMarkers against their classic paper-printed counterparts, a series of comparative experiments were conducted using both marker variants in the ArUco format. The experiments focused on two primary objectives: (1) determining the maximum detection range across varying viewing angles, and (2) assessing the accuracy of the estimated 6DoF pose for each marker type; both were done using the ArUco detection algorithm. The employed sensor across all scenarios was the single-vision setup with Intel® RealSense™ D435, unmodified for printed markers and equipped with static/dynamic polarizers for iMarker detection. We intentionally limited the analysis to this setup to ensure fair comparison for detecting printed markers and iMarkers, while other sensors in Fig. 6 differ significantly in resolution, shutter type, and optical characteristics. All experiments were conducted using the camera’s RGB sensor at a resolution of $1024\times 768$ to maintain uniform input quality across evaluations. Regarding marker dimensions, $7\times 7~\mathrm{cm^{2}}$ visible-range iMarker variants were used to be compared against $6\times 6~\mathrm{cm^{2}}$ and $7\times 7~\mathrm{cm^{2}}$ printed ArUco marker references. Evaluations were performed under normal lighting conditions and in a controlled robotics laboratory equipped with an OptiTrack motion capture system, providing sub-millimeter ground truth accuracy for 6DoF pose tracking. Both the printed fiducial markers and the fabricated iMarkers were affixed to a tripod with an adjustable mounting plate, allowing precise control over their orientation relative to the camera. To ensure accurate ground truth acquisition during the evaluations, omnireflective motion capture landmarks were placed on both the fiducial markers, with their poses continuously tracked using the OptiTrack system. The examined viewing angle ranged from $0^{\circ}$ to $75^{\circ}$, within which the estimated poses of the markers remained reliable w.r.t. the ground truth.
@@
Maximum detection range. According to the evaluation results in Fig. 9, increasing the viewing angle reduces the maximum marker detection range in all scenarios, mainly due to perspective distortion. Additionally, iMarker detection using static and dynamic single-vision setups shows negligible differences in maximum detection ranges, differing between $3~\mathrm{cm}$ to $7~\mathrm{cm}$, with an average of $4.82~\mathrm{cm}$. However, regardless of the detection setup, printed markers achieved consistently higher detection ranges than iMarkers of identical dimensions ($7\times 7~\mathrm{cm^{2}}$). The detection range of $7\times 7~\mathrm{cm^{2}}$ iMarkers closely matches that of $6\times 6~\mathrm{cm^{2}}$ printed ArUco markers, with differences ranging from $4.12~\mathrm{cm}$ to $6.39~\mathrm{cm}$ and averaging $4.24~\mathrm{cm}$ across various angles of view. It should be noted that printed ArUco markers, thanks to their high-contrast black-and-white patterns, offer substantial visual features that facilitate detection. However, iMarkers demonstrate comparable detection performance in real-world experiments, apart from their distinct advantage over classic markers: their ability to seamlessly blend into the environment without disrupting visual aesthetics. This makes iMarkers highly suitable for shorter-range applications where visual discretion is essential, such as indoor navigation.
@@
Table 2: Mean pose estimation error (in meters) of $7\times 7~\mathrm{cm^{2}}$ markers across different viewing angles. Results are reported for “printed markers” and “iMarkers” (detected using static single-vision (SSV) and dynamic single-vision (DSV) setups) of ArUco library. Values in brackets indicate the iMarkers’ relative increase in error compared to the printed markers at each angle.
@@
Pose accuracy. The pose estimation accuracy was assessed by measuring the deviation of the detected marker poses from ground truth data under the mentioned viewing angles. As shown in Table 2, the mean pose estimation error norm across various viewing angles remains consistent among the printed ArUco markers and their iMarker counterparts. In general, the detection error increases as the viewing angle grows from $0^{\circ}$ to $75^{\circ}$ due to increasing perspective distortion of the marker appearance. The relative difference in pose estimation error between printed markers and iMarkers remains within a narrow range of $1.3\%$ to $3.9\%$. This consistency and negligible error differences were expected outcomes of our system design, as the proposed hardware-software configurations are explicitly designed to produce pre-processed binary images revealing the marker patterns prior to the detection and pose estimation stage in the standard ArUco detector library. Thus, the marker detector library can recognize both types of markers by applying built-in preprocessing to either standard RGB inputs of printed markers, or the binary outputs generated by the iMarker detection system.

### v5

@@ 4.2 Dynamic Single-vision Setup
Sensor Design. The second approach offers a single camera equipped with a switchable (dynamic) circular polarizer alternating between left- and right-handed polarization (Fig. 5a), which will be described in detail in a forthcoming paper. In brief, it comprises a fixed circular polarizer paired with a nematic liquid crystal film that functions as a $\lambda/2$-plate in its relaxed (OFF) state and loses its optical functionality when an electric field is applied (ON). A $\lambda/2$-plate reverses circular polarization; therefore, by switching the liquid crystal film between its ON and OFF states (controlled by a function generator synchronized with the camera’s frame rate), the camera sequentially captures frames with right- and left-handed polarization. In an ideal setup, camera frame $f_{t}$ captures the scene with a particular polarization (left- or right-handed), and frame $f_{t+1}$ captures it with the opposite polarization. Thus, the function generator should send signal $\mathbf{S}_{g}$ to trigger the liquid crystal film at time $t$ as:
@@
The liquid crystal cell’s OFF-to-ON switching time is negligible, but the ON-to-OFF relaxation is slower. Thus, intermediate frames with undefined polarization at high frame rates may need to be discarded during iMarker detection. Once optimized for the liquid crystal cell, it ensures that CSR-coated regions in the iMarkers are blocked in one processed frame and visible in the next.
Detection Algorithm. Inspired by the spatial frame processing in the dual-vision setup, this approach employs a frame-level algorithm that uses temporal subtraction to detect the iMarker patterns. Assuming the liquid crystal film operates sufficiently fast w.r.t the video frame rate to produce sequential frames with alternating polarization states (Equation 1), the consecutive (temporal) frame subtraction algorithm to detect iMarkers using a switchable polarizer is shown in Algorithm 2. Here, each frame $f_{t-1}^{\text{~off}}$ and its subsequent frame $f_{t}^{\text{~on}}$, captured with alternating polarization states, are subtracted to produce a grayscale image $f_{p}$, which can then be enhanced through post-processing to facilitate iMarker detection. Additionally, the current frame is stored for subtraction from its succeeding frame in the next iteration. As with the previous approaches, the final stage involves detecting markers $\mathbf{M}$ within the frame $f_{p}$.
@@ 4.3 Static Single-vision Setup
Sensor Design. Another configuration offers a solution for iMarker detection using a single camera equipped with a static circular polarizer attached to its lens (Fig. 5b). The polarizer is selected to block the circular polarization reflected by the CSRs forming the iMarker pattern. For instance, a left-handed circular polarizer will block iMarker regions with CSRs reflecting right-handed polarization, and vice versa. Thus, a visible-range iMarker with CSRs reflecting right-hand polarized light on a same-color background (e.g., green), indistinguishable to the human eye or standard cameras, becomes detectable as the sensor’s left-handed polarizer blocks the CSR regions, revealing the camouflaged iMarker patterns. This design is well suited to detecting visible-range iMarkers camouflaged in the environment, and it can also be used for IR/UV-range iMarkers with appropriate cameras.
Detection Algorithm. Unlike the dual-vision solution, the algorithm for the static single-vision setup does not eliminate the background. However, it simplifies the integration of an iMarker detection sensor by shifting more computation to the detector, achieving a practical balance between complexity and performance. For the visible-range camouflaged iMarkers, the procedure is shown in Algorithm 3. Accordingly, detecting iMarkers $\mathbf{M}$ in a frame $f\in\mathbf{F}$ involves RGB to HSV color space conversion to isolate the color range corresponding to the marker. The resulting mask frame $f_{p}$ highlights the target color range $c_{r}$ as black pixels, with all other areas appearing as white. With the assistance of the polarizer, the mask frame $f_{p}$ highlights the blocked CSR regions of an iMarker with color $c$ (e.g., green) positioned against a background of the same color. Subsequent stages apply post-processing to eliminate noise, followed by detecting marker patterns $\mathbf{M}$ within $f_{p}$. Detecting UV/IR-range iMarkers follows the same stages as the visible-range detection, provided that a UV/IR-range camera is used. However, color masking is unnecessary because the input is grayscale, and range thresholding is used to identify intensity variations in CSR regions, thereby highlighting the iMarker patterns, as introduced in Algorithm 4.
@@ 5.1 Operational Implementation
Sensor. Fig. 6 depicts the different sensor configurations developed for iMarker detection. The dual-vision setup shown in Fig. 6a comprises two ELP 8MP HD camera modules with $75^{\circ}$ FoV facing a Thorlabs polarizing cube beamsplitter with a 50:50 split ratio, all mounted in a compact cage assembly. The other dual-vision setup in Fig. 6b comprises two synchronized U3-3271LE iDS cameras, aligned with a Thorlabs polarizing plate beamsplitter, housed within a custom-designed 3D-printed cage. The static single-vision setup in Fig. 6c uses an Intel® RealSense™ D435 camera with a left-handed circular polarizer affixed to its RGB lens, effectively blocking the right-handed CSR-coated areas of camouflaged iMarkers. Finally, the dynamic single-vision sensor setup in Fig. 6d incorporates a switchable polarizer controlled by a function generator mounted on an iPhone 13 case and aligned with its camera. While this prototype relies on manual polarizer switching, we are investigating a version in which the detection software dynamically synchronizes the camera and the function generator to adjust the polarization.
@@ 5.2 Qualitative Evaluation
Given the complexity of measuring aesthetic perception, this section aims to visually assess iMarkers and showcase their detection performance. To achieve this, various iMarkers were strategically placed in fixed real-world locations and captured by multiple sensors, highlighting their seamless integration across diverse settings. While aesthetic perception is inherently subjective, the visual comparisons in Fig. 1 and Fig. 3 provide practical, real-world evidence supporting the unobtrusiveness of iMarkers across diverse environments.
@@
Fig. 8 depicts several qualitative results, showing the versatility of iMarkers across different spectral ranges and detection setups. Accordingly, Fig. 8(a) shows a green iMarker in the visible range discreetly embedded on a similarly colored background, demonstrating its visual camouflage and unobtrusiveness. In this scenario, the static single-vision setup was employed to capture the scene, with the setup’s polarizer effectively blocking the CSR regions that form the inner iMarker pattern, thereby enhancing contrast and improving color separation for detection. The detection and pose estimation outcome was obtained using Algorithm 3. Fig. 8(b) shows the output of the dynamic single-vision setup applied to a visible-range iMarker under indoor conditions with controlled illumination. In contrast, Fig. 8(c) presents the corresponding outdoor results, where varying sunlight, shadows, reflections, and background clutter introduce additional challenges for reliable iMarker detection. Notably, the system properly isolates the marker region with high contrast, enabling reliable pattern recognition and data decoding. In both dynamic single-vision scenarios, detection is achieved through temporal subtraction between frames $f_{t-1}^{\text{~off}}$ and $f_{t}^{\text{~on}}$ following the procedure described in Algorithm 2. Fig. 8(d) presents the output of the dual-vision setup applied to a visible-range iMarker placed on a patterned surface under low-illumination conditions. In this configuration, both cameras simultaneously capture the same scene, with the left camera acquiring frame $f_{l}$ through a left-handed polarizer and the right camera capturing frame $f_{r}$ through a right-handed polarizer. Marker detection was performed using Algorithm 1, which applies spatial subtraction between the polarized image pairs. The system effectively suppresses background texture and illumination artifacts, preserving high contrast and reliable detection under challenging lighting conditions. As illustrated in Fig. 8(e), the UV-range iMarker placed on a transparent surface is virtually invisible to the naked eye. To reveal the inner patterns of the iMarker, near-UV illumination was applied, and the same scene was captured with a UV-range camera, as shown in the middle sub-figure. The method introduced in Algorithm 4 facilitates marker detection by applying a tailored thresholding strategy to the UV image, enabling reliable extraction of iMarker’s invisible patterns.
Across the evaluated scenarios, iMarkers consistently remain detectable under various environmental and sensing conditions, including indoor/outdoor settings, varying illumination, patterned backgrounds, and transparent surfaces. Despite reduced visual contrast, the proposed sensing and detection strategies reliably isolate the marker regions and preserve sufficient pattern contrast for information decoding.
@@ 5.3 Detection Range and Pose Accuracy Evaluation
To evaluate the effectiveness of iMarkers compared with their paper-printed counterparts, a series of comparative experiments was conducted using both marker variants in the ArUco format. The experiments focused on two primary objectives: (1) determining the maximum detection range across varying viewing angles, and (2) assessing the accuracy of the estimated 6DoF pose for each marker type; both were done using the ArUco detection algorithm. The employed sensor across all scenarios was the single-vision setup with the RealSense, unmodified for printed markers and equipped with static/dynamic polarizers for iMarker detection. We intentionally limited the analysis to this setup to ensure a fair comparison for detecting markers, whereas the other sensors in Fig. 6 differ significantly in resolution, shutter type, and optical characteristics. All experiments were conducted using the camera’s RGB sensor at a resolution of $1024\times 768$ to maintain uniform input quality across evaluations. Regarding marker dimensions, $7~\mathrm{cm}\times 7~\mathrm{cm}$ visible-range iMarker variants were used to be compared against $6~\mathrm{cm}\times 6~\mathrm{cm}$ and $7~\mathrm{cm}\times 7~\mathrm{cm}$ printed ArUco marker references. Evaluations were performed under normal lighting conditions in a controlled robotics laboratory equipped with an OptiTrack motion capture system, providing sub-millimeter ground-truth accuracy for 6DoF pose tracking. Both markers were affixed to a tripod using an adjustable mounting plate, enabling precise control of their orientation relative to the camera. To ensure accurate ground-truth acquisition during the evaluations, omnireflective motion-capture landmarks were placed on both fiducial markers, and their poses were continuously tracked using the OptiTrack system. The examined viewing angle ranged from $0^{\circ}$ to $75^{\circ}$, within which the estimated poses of the markers remained reliable w.r.t. the ground-truth.
@@
Maximum detection range. According to the evaluation results in Fig. 9, increasing the viewing angle reduces the maximum marker detection range in all scenarios, mainly due to perspective distortion. Additionally, iMarker detection using static and dynamic single-vision setups shows negligible differences in maximum detection ranges, differing between $3~\mathrm{cm}$ to $7~\mathrm{cm}$, with an average of $4.82~\mathrm{cm}$. Regardless of the detection setup, printed markers consistently achieved higher detection ranges than iMarkers of identical dimensions ($7~\mathrm{cm}\times 7~\mathrm{cm}$). The detection range of $7~\mathrm{cm}\times 7~\mathrm{cm}$ iMarkers closely matches that of $6~\mathrm{cm}\times 6~\mathrm{cm}$ printed ArUco markers, with differences ranging from $4.12~\mathrm{cm}$ to $6.39~\mathrm{cm}$ and averaging $4.24~\mathrm{cm}$ across various angles of view. It should be noted that printed ArUco markers, thanks to their high-contrast black-and-white patterns, offer substantial visual features that facilitate detection. However, iMarkers demonstrate comparable detection performance in real-world experiments, apart from their distinct advantage over classic markers: their ability to seamlessly blend into the environment without disrupting visual aesthetics. This makes iMarkers highly suitable for shorter-range applications where visual discretion is essential, such as indoor navigation.
@@
Table 2: Mean pose estimation error (in meters) of $7\times 7~\mathrm{cm^{2}}$ markers across different viewing angles. Results are reported for “printed markers” and “iMarkers” (detected using static single-vision (SSV) and dynamic single-vision (DSV) setups) of the ArUco library. Values in brackets indicate the iMarkers’ relative increase in error compared to the printed markers at each angle.
@@
Pose accuracy. The pose estimation accuracy was assessed by measuring the deviation between the detected marker poses and the ground-truth at specified viewing angles. As shown in Table 2, the mean pose estimation error norm across various viewing angles remains consistent among the printed ArUco markers and their iMarker counterparts. In general, the detection error increases as the viewing angle grows from $0^{\circ}$ to $75^{\circ}$due to increasing perspective distortion of the marker appearance. The relative difference in pose estimation error between printed markers and iMarkers remains within a narrow range of $1.3\%$ to $3.9\%$. This consistency and negligible differences in error were expected outcomes of our system design, as the proposed hardware-software configurations are explicitly designed to produce pre-processed binary images that reveal the marker patterns before the detection and pose estimation stage in the standard ArUco detector. Thus, the marker detector library recognizes both markers by applying built-in preprocessing to either standard RGB inputs of printed markers or the binary outputs generated by the iMarker detector.
Results analysis. It should be noted that the goal of iMarkers is not to outperform printed markers in all performance metrics, but to provide an imperceptible yet reliably detectable alternative. Traditional printed markers benefit from high contrast patterns (e.g., black and white) and inherently offer superior detectability than all blended/unobtrusive systems reviewed in Section 2.2. While the invisibility of iMarkers introduces trade-offs in detection robustness, our experimental evaluations of maximum detection range and pose accuracy show that iMarkers can still deliver competitive results, while maintaining minimal visual intrusion.
5.4 Operation in Degraded Visual Conditions
To further highlight the complementary nature of iMarkers relative to conventional markers, an additional evaluation examines scenarios in which iMarkers remain detectable while printed markers fail. These experiments target degraded visual conditions, including low-illumination and near-dark environments, where visible-spectrum contrast is insufficient for reliable detection of printed markers. To assess real-world applicability in robotics, a representative UAV-landing–inspired marker-detection and positioning task was evaluated under varying low-light conditions. This protocol reflects realistic operational scenarios, such as warehouse automation and night-time inspection, in which robots need to operate reliably under severely degraded visible light.
The comparison is conducted between an IR-range iMarker and a printed ArUco marker of identical physical size ($6~\mathrm{cm}\times 6~\mathrm{cm}$), both placed at a distance of $1~\mathrm{m}$ from the camera. Markers use the same ArUco dictionary (original) and differ only in their marker_id to support flexible configurations. Ambient illuminance was gradually reduced during the experiment by controlling the environmental light sources, with illuminance levels measured using a calibrated lux meter application. Furthermore, frames captured at matched resolution using the RGB sensor and the IR-range camera with active IR illumination (Ultrafire A100-IR LED torch) and a long-pass filter, with $18$ synchronized frames sampled per illumination level for both RGB and IR streams across all scenarios. Hence, the RGB stream is used for printed marker detection, and the infrared stream for iMarker detection.
Table 3: Detection robustness of printed ArUco markers ($\mathcal{P}$) and IR-range iMarkers ($\mathcal{I}$) under progressively lower illumination conditions, highlighting scenarios where conventional markers fail but iMarkers remain reliably detectable.
Illuminance
$15~\mathrm{lux}$ $10~\mathrm{lux}$ $5~\mathrm{lux}$ $\leq 5~\mathrm{lux}$
Sample View $\mathcal{P}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\mathcal{I}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Sample Detection $\mathcal{P}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\mathcal{I}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Detection Rate $\mathcal{P}$ 0.72 (13 / 18) 0.50 (9 / 18) 0.11 (2 / 18) 0.06 (1 / 18)
$\mathcal{I}$ 0.89 (16 / 18) 0.83 (15 / 18) 0.83 (15 / 18) 0.78 (14 / 18)
# False Positive $\mathcal{P}$ 0.17 (3 / 18) 0.33 (6 / 18) 0.33 (6 / 18) 0.44 (8 / 18)
$\mathcal{I}$ 0.06 (1 / 18) 0.11 (2 / 18) 0.17 (3 / 18) 0.17 (3 / 18)
As summarized in Table 3, various complementary metrics, including detection rate and false positive frequency, have been analyzed for printed markers and IR-range iMarkers under progressively decreasing illumination conditions. Here, the detection rate is defined as the ratio of frames in which the correct marker_id is detected to the total number of frames captured, all at illumination level $l$. False positives are incorrect detections triggered by high-contrast background textures that resemble marker patterns, such as repetitive grids or net-like structures.
According to the table, the detection rate of printed markers degrades sharply as illumination decreases, because reduced contrast causes marker features to blend into background noise. Specifically, performance drops from $0.72$ at 15lux15~lux to 0.06 at $\leq 5~\mathrm{lux}$. In contrast, IR iMarkers maintain a largely stable detection rate, varying only slightly from 0.89 to 0.78, with minor degradation attributed to sensor noise and IR signal attenuation under extremely low-light conditions. A similar trend is observed for false positives: as illumination decreases, erroneous detections for printed markers increase substantially (from 0.17 to 0.44), primarily due to the reduced signal-to-noise ratio in the RGB stream. In contrast, iMarkers maintain a consistently low false-positive rate across all illumination levels, with only minor fluctuations, owing to the CSRs’ distinctive IR reflectance properties.
These results show that while printed markers are strongly affected by illumination-dependent failure modes, iMarkers provide reliable detection performance under degraded visual conditions. Additionally, it clearly demonstrates that iMarkers go beyond aesthetic considerations and act as an enabling technology for robotic perception in scenarios where conventional fiducial markers become unreliable or unusable.